The length of one leg of a right triangle is 15 feet. The length of the hypotenuse is 9 feet longer than the other leg.
Find the length of the hypotenuse and the other side.

Answer by stanbon(57361)

The length of one leg of a right triangle is 15 feet. The length of the hypotenuse is 9 feet longer than the other leg.
Find the length of the hypotenuse and the other side.
--------------------
Let the other leg be "x".
Then hypotenuse = "x+9".
----
Equation:
leg^2 + otherleg^2 = hypotenuse^2
15^2 + x^2 = (x+9)^2
225 + x^2 = x^2+18x+81
---
18x = 144
x = 8 (leg)
x+9 = 17 (other leg)
=======================
Cheers,
Stan H.
=================